Equations with two variables.
-Find the Y value for the equation for x = -2, 0, and 2.
y=3x-4
Thank you

Answer by Leaf W.(135)

Plug in each value for x and simplify--
For x = -2 :
y = 3(-2) - 4
y = -6 - 4
y = -10
For x = 0 :
y = 3(0) - 4
y = 0 - 4
y = -4
For x = 2 :
y = 3(2) - 4
y = 6 - 4
y = 2
Therefore, y = -10 when x = -2, y = -4 when x = 0, and y = 2 when x = 2.